OpenStudy (anonymous):

nitrogen gas occupies 6.5L at 27 C and 5 atm pressure. What will the volume be at -100 C and the same pressure

OpenStudy (anonymous):

ok for this kind of problems you should always use pV=nRT and in this case you have same amount of gass but different variables so: n1=n2 (1and 2 are states of gas at given pressure, temperature and volume) n=pV/RT p1V1/T1=p2V2/T2 cause R is constant it strikes out, now you have that pressure is the same so: V1/T1=V2/T2 so V2 or volume at different temperature is: V2= V1 T2/T1

OpenStudy (unklerhaukus):

and don't forget to convert those temperatures to Kelvin